OpenStudy (anonymous):

how do i convert to rectangular coordinates? rcsc(Theta) = -2

OpenStudy (tkhunny):

\(r\csc(\theta) = \dfrac{r}{\sin(\theta)} = \dfrac{r^{2}}{r\sin(\theta)}\)